OpenStudy (anonymous):

the product of two positive integers is 98. if they differ by 7, what are the integers?

OpenStudy (anonymous):

xy=98 x-y=7 solving these gives x=14, y=7

OpenStudy (anonymous):

x*y = 98 x - y = 7 x = 14, y = 7